# Scalable Valuation of Human Feedback through Provably Robust Model Alignment

### v1

@@ 5.1 Controlled Sentiment Generation: Robustness Evaluation
Setup. We evaluate alignment robustness on a sentiment-controlled text generation task using the IMDb dataset [68], which contains 25,000 polarized movie reviews. Following Chowdhury et al. [23], we extract the first 20 tokens from each review as a context prompt. For each prompt, we generate four responses using the SFT-ed GPT2-large [81]. These responses are used to construct six preference pairs per prompt. To provide sentiment rankings, we employ SiEBERT [66, 45] as the latent (ground-truth) reward model $r^{*}(y,x)$. To ensure clean dataset, we filter out preference pairs that do not satisfy $r^{*}(y_{\text{win}},x)-r^{*}(y_{\text{lose}},x)>0.1$, yielding 12,000 clean pairs. Of these, 10,000 are used for training and 2,000 for evaluation. To simulate label noise, we randomly flipped $(y_{\text{win}},y_{\text{lose}})$ in the training set with contamination ratio $\epsilon\in\{0.1,0.2,0.3,0.4\}$. We then fine-tuned GPT2-large using various DPO variants. Alignment quality is measured by the average reward assigned by SiEBERT to responses generated by the aligned model on the 2,000 evaluation prompts.
Alignment quality. Figure 3(a) shows that Hölder-DPO consistently outperforms all baselines in average reward, with only a minor drop as the contamination level $\epsilon$ increases. This supports our theoretical claim: Hölder-DPO, uniquely characterized by the redescending property, achieves robustness even under severe noise (i.e., when $\epsilon\geq 0.3$). A mild linear decay remains, likely due to the reduced proportion of effectively clean data points, which limits estimation accuracy. Figures 3(b) and 3(c) further demonstrate Hölder-DPO’s robustness across generation temperatures and training steps. Notably, while the baseline methods plateau early during training, Hölder-DPO continues to improve, suggesting stronger resistance to overfitting on noisy preferences.
Valuation quality. Figures 3(d) and 3(e) evaluate mislabel valuation. As shown in Proposition 1, this estimation capability is unique to Hölder-DPO and is not supported by other baselines (see Theorem 3). The detection task is entirely unsupervised—no ground-truth labels indicate whether a preference pair is clean or flipped. Remarkably, Hölder-DPO accurately estimates the contamination ratio up to $\epsilon=0.2$, after which the estimates begin to saturate. This is expected, given that contamination ratio estimation relies on the optimized $\sigma(g_{\theta^{*}(\epsilon)})$, and the average reward slightly deteriorates as $\epsilon$ increases (see Figure 3(a)). Nonetheless, it reliably ranks datasets by relative noise levels. For mislabel detection, precision—i.e., the proportion of true mislabels among those predicted as mislabeled—increases with higher $\epsilon$, which we attribute to underestimation of $\epsilon$, reducing false positives and improving detection precision.
Impact of $\gamma$ choice.
@@
$\gamma$ average reward $\hat{\epsilon}$ precision
@@
We vary the key hyperparameter $\gamma$ to assess its effect. As shown in Table 1, increasing $\gamma$ improves contamination ratio estimation $\hat{\epsilon}$ but lowers mislabel detection precision, revealing a trade-off. Intuitively, $\gamma$ controls the level of contamination removal. While tuning $\gamma$ to match the true $\epsilon$ yields optimal performance, the true $\epsilon$ is typically unknown. Importantly, even with a fixed $\gamma=2$ across all experiments, Hölder-DPO consistently outperforms all baselines, suggesting that extensive hyperparameter tuning is unnecessary. See Appendix H for additional ablations on other hyperparameters, including batch size.
@@ 5.2 Single-turn Dialogue: Generated Responce Quality Evaluation
Figure 4: Helpful assistant dialogue generation on the Golden HH dataset. Error bars denote standard deviation over 10 trials with different random seeds. Hölder-DPO consistently achieves the highest GPT-4 win rate across both base models: (a)Qwen-2.5-1.5B and (b)Phi-2. Notably, Hölder-DPO is the only method that outperforms GPT-4-generated prompts even when using the smaller 2.8B Phi-2 model. It also delivers near-perfect (d) contamination ratio estimation $\hat{\epsilon}$ and (e) precision in mislabeled data detection, regardless of the base model.
Given a user prompt, the task is to generate responses that are more helpful and less harmful. In this setting, we do not have access to ground-truth rewards (e.g., from a sentiment model like SiEBERT). Instead, we adopt the GPT-4 win rate [82, 103], the standard metric for evaluating alignment when ground truth is unavailable. This metric follows three steps: (A) generate a response $y_{\text{gen}}$ using the aligned model for a given prompt $x$; (B) retrieve the preferred response $y_{\text{win}}$ from the test dataset for the same prompt $x$; (C) ask GPT-4 to compare $(y_{\text{win}},y_{\text{gen}})$ and determine which is better. The win rate is defined as the fraction of prompts where $y_{\text{gen}}$ is preferred. For dataset, we evaluate on the Golden HH [15], a manually curated version of Anthropic HH [33], in which low-quality responses have been replaced with GPT-4-generated outputs. Since $y_{\text{win}}$ corresponds to GPT-4 generations, the benchmark is both cleaner and more challenging than the original Anthropic HH. The dataset contains 42,500 training and 2,310 test examples. To evaluate robustness, we randomly flipped preference labels with varying noise levels $\epsilon\in\{0.,0.1,0.2,0.3,0.4\}$. We tested two base models: Qwen-2.5-1.5B [105] and Phi-2 [49], and fine-tuned using DeepSpeed Stage 3 [3] for memory efficiency, with an effective batch size of 32222Batch size of 4 with gradient accumulation of 8.. As shown in Figures 4(a) and 4(b), Hölder-DPO consistently outperforms all baselines across both base models. Notably, Hölder-DPO with Phi-2 is the only method to exceed a 50% win rate, slightly outperforming GPT-4-generated responses. Figures 4(c) and 4(d) further show that mislabel detection is nearly perfect. Remarkably, these results are achieved in a fully unsupervised setting, underscoring the effectiveness of our approach.
@@
5.3 Dataset Valuation and Cleaning
@@
We next applied Hölder-DPO to real-world noisy datasets for dataset valuation and cleaning via our contamination ratio estimator $\hat{\epsilon}$ in Eq. (11), where the true contamination ratio $\epsilon$ is unknown. Specifically, we used the Anthropic HH dataset [33], which is widely used but known to contain label noise [86, 19, 103]. We adopted Phi-2 as the base model and apply Hölder-DPO directly to the raw dataset.
@@
Data cleaning. We also evaluated the effectiveness of Hölder-DPO for data cleaning. Here, Hölder-DPO was used as a pre-processing step to identify and remove noisy preference labels, yielding a filtered training dataset. We then compared the GPT-4 win rates before and after cleaning. As shown in Figure 5(c), removing samples identified as mislabeled consistently improves performance across all methods. Notably, both DPO and Dr. DPO trained on the cleaned dataset outperform Hölder-DPO trained on the original (noisy) dataset. This is expected: while Hölder-DPO is designed to be robust under label noise ($\epsilon>0$), once the data is cleaned ($\epsilon\approx 0$), simpler methods like DPO can align more effectively. These results suggest that Hölder-DPO can serve not only as a robust training objective, but also as a valuable pre-filtering tool for noisy preference datasets.
@@ 6 Conclusion and Limitation
Conclusion. We analyzed the robustness of DPO variants through the lens of the redescending property and showed that none of the existing methods satisfy it. We then introduced Hölder-DPO—the first algorithm with a provable redescending property—alongside a principled data valuation method for estimating contamination and identifying mislabeled data. Our experiments demonstrate that Hölder-DPO improves alignment quality and accurately detects mislabels without access to a clean validation set. Notably, it offers the first theoretically grounded approach to dataset valuation, revealing significant noise in the widely used Anthropic HH dataset. Moreover, filtering out the detected noisy examples further improves performance across alignment methods, positioning Hölder-DPO as an effective pre-training filter. Taken together, our theoretical and empirical results advance scalable, automatable alignment without relying on clean validation data.
Limitations. Hölder-DPO introduces an additional hyperparameter, $\gamma$. While our experiments show robust performance with a default value of 2.0, the optimal choice may vary across tasks and datasets. This sensitivity may require tuning to fully realize Hölder-DPO’s potential. Future work could explore adaptive hyperparameter selection [104] to automate $\gamma$ tuning. Another limitation lies in the coarse approximation of the DP divergence (see Appendix D); exploring more precise approximation methods is an important direction for future work.

### v2

@@ 5.1 Controlled Sentiment Generation: Robustness Evaluation
Setup. We evaluate alignment robustness on a sentiment-controlled text generation task using the IMDb dataset [71], which contains 25,000 polarised movie reviews. Following Chowdhury et al. [24], we extract the first 20 tokens from each review as a context prompt. For each prompt, we generate four responses using the SFT-ed GPT2-large [83]. These responses are used to construct six preference pairs per prompt. To provide sentiment rankings, we employ SiEBERT [69, 46] as the latent (ground-truth) reward model $r^{*}(y,x)$. To ensure clean dataset, we filter out preference pairs that do not satisfy $r^{*}(y_{\text{win}},x)-r^{*}(y_{\text{lose}},x)>0.1$, yielding 12,000 clean pairs. Of these, 10,000 are used for training and 2,000 for evaluation. To simulate label noise, we randomly flipped $(y_{\text{win}},y_{\text{lose}})$ in the training set with contamination ratio $\epsilon\in\{0.,0.1,0.2,0.3,0.4\}$. We then fine-tuned GPT2-large using various DPO variants. Alignment quality is measured by the average reward assigned by SiEBERT to responses generated by the aligned model on the 2,000 evaluation prompts.
Alignment quality. Figure 3(a) shows that Hölder-DPO consistently outperforms all baselines in average reward, with only a minor drop as the contamination level $\epsilon$ increases. This supports our theoretical claim: Hölder-DPO, uniquely characterised by the redescending property, achieves robustness even under severe noise (e.g., when $\epsilon\geq 0.3$). A mild linear decay remains, likely due to the reduced proportion of effectively clean data points, which limits estimation accuracy. Figures 3(b) and 3(c) further demonstrate Hölder-DPO’s robustness across generation temperatures and training steps. Notably, while the baseline methods plateau early during training, Hölder-DPO continues to improve, suggesting stronger resistance to overfitting on noisy preferences.
@@
$\gamma$
average
reward
$\hat{\epsilon}$ precision
@@
Valuation quality. Figures 3(d) and 3(e) evaluate mislabel valuation. As shown in Proposition 1, this estimation capability is unique to Hölder-DPO and is not supported by other baselines (see Theorem 3). The detection task is entirely unsupervised—no ground-truth labels indicate whether a preference pair is clean or flipped. Remarkably, Hölder-DPO accurately estimates the contamination ratio up to $\epsilon=0.2$, after which the estimates begin to saturate. This is expected, given that contamination ratio estimation relies on the optimised $\sigma(g_{\theta^{*}(\epsilon)})$, and the average reward slightly deteriorates as $\epsilon$ increases (see Figure 3(a)). Nonetheless, it reliably ranks datasets by relative noise levels. For mislabel detection, precision—i.e., the proportion of true mislabels among those predicted as mislabelled—increases with higher $\epsilon$, which we attribute to underestimation of $\epsilon$, reducing false positives and improving detection precision.
Impact of $\gamma$ choice.
We vary the key hyperparameter $\gamma$ to assess its effect. As shown in Table 1, there is a trade-off: increasing $\gamma$ leads to a more accurate contamination ratio estimate $\hat{\epsilon}$ but lowers mislabel detection precision. Intuitively, $\gamma$ controls how strongly potential outliers are down-weighted. While the optimal $\gamma$ is task-dependent and relates to the true (yet unknown) $\epsilon$, Hölder-DPO consistently outperforms all baselines with a fixed default of $\gamma=2$, suggesting that it is effective without extensive tuning. See Appendix H for further ablation studies on hyperparameters, e.g., batch size.
@@ 5.2 Single-turn Dialogue: Generated Responce Quality Evaluation
Figure 4: Helpful assistant dialogue generation on the Golden HH dataset. Error bars denote standard deviation over 10 trials with different random seeds. Hölder-DPO consistently achieves the highest GPT-4 win rate across both base models: (a)Qwen-2.5-1.5B and (b)Phi-2. Notably, Hölder-DPO is the only method that outperforms GPT-4-generated prompts even when using the smaller 2.8B Phi-2 model. It also delivers near-perfect (d) contamination ratio estimation $\hat{\epsilon}$ and (e) precision in mislabelled data detection, regardless of the base model.
Given a user prompt, the task is to generate responses that are more helpful and less harmful. In this setting, we do not have access to ground-truth rewards (e.g., from a sentiment model like SiEBERT). Instead, we adopt the GPT-4 win rate [84, 105], the standard metric for evaluating alignment when ground truth is unavailable. This metric follows three steps: (A) generate a response $y_{\text{gen}}$ using the aligned model for a given prompt $x$; (B) retrieve the preferred response $y_{\text{win}}$ from the test dataset for the same prompt $x$; (C) ask GPT-4 to compare $(y_{\text{win}},y_{\text{gen}})$ and determine which is better. The win rate is defined as the fraction of prompts where $y_{\text{gen}}$ is preferred. For dataset, we evaluate on the Golden HH [15], a manually curated version of Anthropic HH [34], in which low-quality responses have been replaced with GPT-4-generated outputs. Since $y_{\text{win}}$ corresponds to GPT-4 generations, the benchmark is both cleaner and more challenging than the original Anthropic HH. The dataset contains 42,500 training and 2,310 test examples. To evaluate robustness, we randomly flipped preference labels with varying noise levels $\epsilon\in\{0.,0.1,0.2,0.3,0.4\}$. We tested two base models: Qwen-2.5-1.5B [106] and Phi-2 [50], and fine-tuned using DeepSpeed Stage 3 [3] for memory efficiency, with an effective batch size of 32555Batch size of 4 with gradient accumulation of 8.. As shown in Figures 4(a) and 4(b), Hölder-DPO consistently outperforms all baselines across both base models. Notably, Hölder-DPO with Phi-2 is the only method to exceed a 50% win rate, slightly outperforming GPT-4-generated responses. Figures 4(c) and 4(d) further show that mislabel detection is nearly perfect. Remarkably, these results are achieved in a fully unsupervised setting, underscoring the effectiveness of our approach.
@@
5.3 Real-world Noisy Datasets: Dataset Valuation and Cleaning
@@
To go beyond controlled settings with synthetic noise, we next applied Hölder-DPO to real-world datasets where the true contamination ratio $\epsilon$ is unknown. Using our contamination ratio estimator $\hat{\epsilon}$ in Eq. (11), we performed dataset valuation and cleaning on the Anthropic HH dataset [34], which is widely used but known to contain significant label noise [88, 20, 105]. We adopted Phi-2 as the base model and apply Hölder-DPO directly to the raw dataset.
@@
Data cleaning. We also evaluated the effectiveness of Hölder-DPO for data cleaning. Here, Hölder-DPO was used as a pre-processing step to identify and remove noisy preference labels, yielding a filtered training dataset. We then compared the GPT-4 win rates before and after cleaning. As shown in Figure 5(c), removing samples identified as mislabelled consistently improves performance across all methods. Notably, both DPO and Dr. DPO trained on the cleaned dataset outperform Hölder-DPO trained on the original (noisy) dataset. This is expected: while Hölder-DPO is designed to be robust under label noise ($\epsilon>0$), once the data is cleaned ($\epsilon\approx 0$), robust objectives can be overly conservative, thus KL-based methods like DPO can align more effectively. These results suggest that Hölder-DPO can serve not only as a robust training objective, but also as a valuable pre-filtering tool for noisy preference datasets.
6 Discussion
Role of Hölder-DPO; pre-filtering or new objective?
Hölder-DPO is a versatile framework whose optimal role depends on the application context. When maximising final performance is prioritised over training cost, a two-stage pipeline—first filtering with Hölder-DPO and then fine-tuning using a vanilla DPO objective—yields superior results (see Figure 5(c)). Conversely, when computational efficiency is the priority, a single-stage approach—directly fine-tuning with the Hölder-DPO objective—offers a cost-effective solution while maintaining strong alignment quality. This flexibility enables practitioners to choose the two strategies according to their constraints.
How to set $\gamma$?
As shown in Table 1, the value of $\gamma$ has a strong influence on performance. Its optimal setting is likely task- and dataset-dependent. Ideally, dynamically tuning $\gamma$ during training would yield the best results. Nevertheless, our experiments show that a default value of 2.0 performs robustly across diverse settings, and we recommend using it as the default choice. When fine-tuning $\gamma$ is desired, a promising approach is to employ a human-in-the-loop process. As illustrated in Figure 5(b), the Hölder-DPO model can rank data points. This ranking can serve as a diagnostic tool to guide manual inspection: for instance, increasing $\gamma$ if clear mislabels are still being tolerated, or decreasing it if correct preferences are penalised as mislabels.
What if $\epsilon>0.5$?
When the noise ratio becomes extreme (e.g., $\epsilon=0.9$), Hölder-DPO can still identify the 10% of clean samples by isolating them as outliers. In such cases, the roles can be reversed—the identified minority can be treated as the clean subset, while the majority is discarded. This reversal can be validated with minimal human effort by inspecting a few of the top-ranked samples flagged by our method. Hence, Hölder-DPO remains applicable even in highly noisy settings (e.g., $\epsilon>0.6$). However, detecting mislabeled data becomes fundamentally difficult when $\epsilon\approx 0.5$. We argue that this challenge is not a limitation of our method per se, but rather a theoretical identifiability barrier: when clean and noisy data are equally represented, distinguishing between them is statistically impossible without access to a clean reference set. Our approach inherently assumes a bimodal distribution, where one mode—typically the majority—corresponds to clean data.
What kind of label noise does the tail assumption cover?
While our tail-contamination assumption formally presumes i.i.d. noise, it is in fact more general than it may appear. Hölder-DPO is effective against a wide range of label noise types—symmetric, asymmetric, and even certain systematic annotator biases—so long as the resulting noisy preference pairs are recognized by the learned model as low-likelihood events (i.e., outliers). This generality represents a clear advantage over prior methods that can only handle symmetric label flips (e.g., [67]). The main limitation arises when the noise is too structured to be identified as an outlier by the model. For instance, if all preference labels associated with a single prompt or narrow topic are consistently flipped, such errors cease to form a distinct “tail” in the data distribution. Instead, they produce a coherent but incorrect signal within that subdomain—one the model may interpret as a genuine, domain-specific preference. Addressing such structured noise requires modeling dependencies across samples rather than treating each observation as independently corrupted. Exploring this direction is a key avenue for future work. Concurrent research has begun tackling these challenges using latent-variable formulations [16].
@@ 7 Conclusion and Limitation
Conclusion. We analyzed the robustness of DPO variants through the lens of the redescending property and showed that none of the existing methods satisfy it. We then introduced Hölder-DPO—the first algorithm with a provable redescending property—alongside a principled data valuation method for estimating contamination and identifying mislabelled data. Our experiments demonstrate that Hölder-DPO improves alignment quality and accurately detects mislabels without access to a clean validation set. Notably, it offers the first theoretically grounded approach to dataset valuation, revealing significant noise in the widely used Anthropic HH dataset. Moreover, filtering out the detected noisy examples further improves performance across alignment methods, positioning Hölder-DPO as an effective pre-training filter. Taken together, our theoretical and empirical results advance scalable, automatable alignment without relying on clean validation data.
Limitations. A key limitation lies in the selection of the hyperparameter $\gamma$ and see Section 6 for the potential solutions. Another limitation is that our theoretical guarantees are also based on an i.i.d. label flip model and see also Section 6 for the details. Furthermore, our objective relies on a coarse approximation of the DP divergence (see Appendix D), and exploring more precise approximation methods is another important avenue. Finally, we implicitly assume that majority opinion represents the clean data; see Appendix A for the broader implications.